OpenStudy (anonymous):

Compute the following integral. Clearly show the substitution used for the following integral and how it was used. (8y+15/36+y^2)dy sorry i couldnt get that sign for integral that should be at the front.

OpenStudy (mimi_x3):

Break it up: \[\int\limits\frac{8y}{36+y^{2}} dy +\int\limits\frac{15}{36+y^{2}} dy\] Then integrate it one by one. The first one is an inverse function. And the second is a u-substitution.

sam (.sam.):

\[\int\limits_{}^{}\frac{8y}{36+y ^{2}}dy=4\ln(36+y ^{2})+C\] First one

OpenStudy (mimi_x3):

wait..they are both trig inverse functions..my bad.

OpenStudy (anonymous):

yes second function is an arctan

sam (.sam.):

Second one \[\frac{\tan ^{-1}(\frac{y}{6})}{6}+C\]

OpenStudy (diyadiya):

\[\int\limits\limits\frac{15}{36+y^{2}} dy = 15\int\limits\limits\frac{1}{6^2+y^{2}} dy = \frac{15}{6} \tan^{-1} \frac{y}{6}= \frac{5}{2} \tan^{-1} \frac{y}{6}\] So finally \[4\ln(36+y ^{2}) +\frac{5}{2} \tan^{-1} \frac{y}{6} +C\]

OpenStudy (diyadiya):

No here a is 6

OpenStudy (diyadiya):

a^2+ y^2 = 36 + y^2 = (6)^2 + y^2

OpenStudy (mimi_x3):

\[\int\limits\frac{1}{a^{2}+x^{2}} dx = \frac{1}{a} \tan^{-1}\frac{x}{a} \]

OpenStudy (diyadiya):

\[\frac{1}{6} \tan^{-1} \frac{y}{6}\]

OpenStudy (mimi_x3):

Diya, you were correct at the first place..

OpenStudy (diyadiya):

15 \[ 15\int\limits\limits\limits\frac{1}{6^2+y^{2}} dy = 15 \times \frac{1}{6} \tan^{-1} \frac{y}{6}= \frac{5}{2} \tan^{-1} \frac{y}{6} \]

OpenStudy (anonymous):

Thank you guys. I hope i get ready for my exam. have to keep trying to understand though.

OpenStudy (diyadiya):

Goodluck :)

OpenStudy (anonymous):

Thanks